OpenStudy (anonymous):

18. For which system of equations would elimination be the simplest technique to choose? (Points : 2) y = 2x and x + 3y = 5 4x + 3y = 6 and 3y = 5 x + y = 7 and x - y = 12 3x - 2y = 8 and 4x - 5y = 20 19. Substitution would be the simplest method of solution for which system? (Points : 2) 5/3 x = 6y and 2x -3y = 8 y = 2x and x + y = 28 3x - 2y =18 and 4x + 5y =60 20. The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? (Points : 2) 56 41

OpenStudy (anonymous):

18. 4x + 3y = 6 and 3y = 5 because 3y=5 can be sub. into the first equation directly 19. y = 2x and x + y = 28 because y=2x can be sub. into x+y=28 directly 20. x+y=110 x-y=28 By solving these two equations, x=69, y=41 so, 41 is the smallest one

OpenStudy (anonymous):

Oh thank you so much CoCO Tsoi :)

OpenStudy (anonymous):

You are welcome :D